# Active Label Correction for Building LLM-based Modular AI Systems

### v2

@@ 3.2 Iterative Data Correction
The flagged examples obtained from the misannotation predictor are fed to a human annotator for correction. Once corrected by a human expert, these examples can be used to update the dataset. We can perform another iteration of MP including the model training on the updated dataset to obtain a new set of flagged examples. This process can be repeated to iteratively improve the quality of the data until a stopping criterion is met. Stopping conditions will be discussed in the results section. For DALC, in each iteration, we revert the previously auto-flipped examples and auto-flip a new set examples based on new model predictions. Further, an example corrected by human annotation cannot be auto-flipped or flagged in subsequent iterations.
With each iteration, new information obtained from the human annotator is accommodated into the dataset which is used to train the next model for MP. The new model predictions reveal new inconsistencies with the noisy annotations and lead to the next set of flagged examples. As the count of misannotated examples reduces in the noisy dataset with each iteration, we can expect the MP performance to drop. We discuss this phenomenon in the results section. One way to control the MP performance is to control $M$, the fraction of flagged examples. Ideally, one can train a new model after correcting every single example but this extreme case is expensive and infeasible due to the long wait time as MP runs after every example is verified. But we still may want to adjust the value $M$ based on a target precision value. Therefore, in practice, the MP precision can be measured for the latest batch of examples and $M$ can be decreased if precision is low or vice versa. Increasing precision by lowering $M$ comes at the cost of low recall and running more iterations. In our experiments, we use a fixed value of $M$ i.e. a fixed number of examples are flagged in each iteration.
@@ 4.1 Performance on Evaluation Tasks
We evaluate the ALC-based pipeline on three different tasks: (1) ATIS (Airline Travel Information System) Intent Classification Hemphill et al. (1990), (2) CoNLL 2003 Named Entity Recognition or NER Tjong et al. (2003) and (3) Question-answering Natural Language Inference or QNLI Wang et al. (2018). The three tasks are common to many conversational AI agents Lin et al. (2023). Intent classification is used for the downstream processing of incoming requests, NER is used for slot-filling (such as train booking) to fill various parts of the input before a database query, and NLI or textual entailment is used in conversational AI agents to confirm the validity of the output in answering the user query. All three datasets are publicly available and span different levels of complexity and sizes as shown in Table 1. Furthermore, the ground truth annotations allow us to compare our pipeline to the oracle performance.
For each task, we created a prompt using the OpenAI Playground and then used the OpenAI API to get outputs for both the train and test split of each dataset using ‘gpt-3.5-turbo’ model with zero temperature i.e. no randomness in output generation. For fine-tuning, we used ‘distilbert-base-uncased’, ‘roberta-base’ and ‘albert-large-v2’ pre-trained transformer models from HuggingFace for ATIS, CoNLL and QNLI tasks respectively based on their public leaderboards [^2] https://paperswithcode.com/ and availability on HuggingFace. The results for each task are shown in the last column FT in Table 1. Model training details are provided in Appendix B.
@@
Table 1: Performance of ChatGPT on train and test set of three tasks: (1) ATIS Intent Classification, (2) CoNLL 2003 Name Entity Recognition, and (3) QNLI Natural Language Inference along with performance of fine-tuned models in FT column.
ATIS dataset consists of 17 unique intents (‘airport’, ‘aircraft’, ‘distance’, etc.) assigned to user queries. A description of all 17 intents was provided to ChatGPT for classification. 40 examples with multiple labels were removed from the dataset to reduce the task to single-label classification. For the CoNLL dataset, an example of entity recognition output was provided in the prompt to describe the format and four types of entities i.e. person, location, organization, and miscellaneous. QNLI dataset consists of pairs of questions and sentences based on the Stanford Question Answering Dataset Wang et al. (2018). We provide the inputs in the prompt and ask if the given passage can be used to answer the question. Prompts and output parsing details are provided in Appendix C.
The performance of ChatGPT on the three tasks are documented in Table 1. Overall, the results agree with Qin et al. (2023) in terms of the difference in performance between a fine-tuned model and zero-shot learning using GPT-3. We observed a large improvement in F1 Score on CoNLL 2003, 68.2% compared to F1 Score of 53.7% in Qin et al. (2023), possibly because of our different prompt instruction which explicitly asks for the classification of each word instead of extracting a list of named entities. We note that ChatGPT has 10-30% lower performance than fine-tuned models across the three tasks. While zero-shot learners such as ChatGPT generalize well to new tasks without training, this result suggests that there is a large performance gap that needs to be bridged when bootstrapping modular AI agents with ChatGPT. Therefore, it is important to develop a method to improve each module and the overall system over time after deploying the LLM-based AI systems.
@@
ChatGPT 84.0% 0.0949
Table 2: Comparison of different noise types, including no noise (None), Random, Label-conditional, Input-conditional, and ChatGPT annotations, on model Accuracy and Class Imbalance for ATIS dataset.
4.2 Noise Characteristics
We compare different types of noise models to understand the noise induced by ChatGPT while annotating the ATIS dataset. We chose the ATIS dataset for this experiment as it has the highest number of classes which can lead to more pronounced noise patterns. We add different types of synthetic noise to the original dataset with the same proportion of misannotations as ChatGPT. Table 2 shows the accuracy of models trained with different noisy versions of ATIS datasets and the class imbalance measured by the KL divergence of label distribution from the original label distribution. Details about noise models are provided in Appendix E.
We find that ChatGPT adds the most detrimental noise while having a similar class imbalance as other noise models. We believe that ChatGPT provides very reasonable false annotations that are inherently hard to detect or ignore which leads to poor test performance. It is also interesting to note that a model trained on ChatGPT-annotated training data performs much better than ChatGPT on the eval dataset (84.0% versus 75.6%). This is likely due to the input data distribution learned by the model leading to better input representations as compared to a generic pre-trained model. This motivates the fine-tuning of task-specific models for applications in modular AI systems as well as the use of active label correction to improve data quality which can lead to further improvement in the performance of modular AI agents.
Figure 2: A 2D projection of ATIS text embeddings for a subset of 7 classes. ChatGPT annotations are indicated by colors while large dots indicate errors.
In Figure 2, we show 2D t-SNE projections Maaten and Hinton (2008) of Contriever text embeddings Izacard et al. (2022) from ATIS dataset. We visualize examples labeled with 7 classes shown in the legend and larger dots indicate the examples misannotated by ChatGPT. We observe that most misannotated examples lie near intersections or between class clusters. This indicates that noise induced by ChatGPT is indeed complex and depends on the text, its true label, and other possible labels. This also conforms with Table 2 where we find random and label-conditional noise to be non-representative models of ChatGPT annotations.
@@ 4.3 Misannotation Prediction
In Figure 3, we compare RLC, ALC, and DALC described earlier for MP on all three datasets annotated using ChatGPT. We use $\delta=0.75,0.98,0.90$ for ATIS, CoNLL and QLNI to keep the proportion of auto-flipped examples between 3% to 6%. We note from Table 1 that 29.8% of the ATIS and 15.1% of the QNLI train dataset were misannotated by ChatGPT. For CoNLL, 57.4% of examples contain errors in token classes though only 9.7% of tokens in the dataset are misclassified.
Figure 3: Misannotation Prediction (MP) precision and recall for ATIS, CoNLL, and QNLI with change in $M$.
We observe that precision goes down for ALC and DALC as $M$ increases for each task. This is because the examples with the highest probability of being misannotated based on $m(.)$ values indeed have a higher rate of being misannotated. In other words, starting from the examples with the highest $m(.)$ values, the human annotator will see a decrease in the MP quality over time as they start verifying examples with lower $m(.)$ values. We also see an increase in recall as $M$ increases which is expected because more misannotated examples are flagged at higher values of $M$. We also note that DALC has consistently lower precision and recall than ALC. The auto-flipped examples in DALC typically have high $m(.)$ as $p_{\theta}(y\neq y^{*}|x)$ is low but these examples cannot be flagged. ALC, on the other hand, flags the examples with the highest $m(.)$ values leading to higher precision and recall. For RLC, the precision is constant and low while recall increases perfectly linearly with $M$ because examples are flagged at random.
For the human annotator, examples flagged by DALC are more likely to be correct as compared to those flagged by ALC. It provides a more balanced set of flagged examples for the correction task which may also avoid the problem of over-reliance of human annotators on model predictions. Further, the overall quality of the data can be higher with DALC because of two sources of true labels viz. auto-flipping and human corrections. Given a specific method, we wish to maximize precision to correct a higher proportion of misannotated examples in each iteration. Therefore, we use low values of $M$ in each step in iterative data correction.
@@ 4.4 Effect of Data Size on MP
We wish to understand how the quantity of data collected by modules within an AI system will impact MP performance. Towards this, we vary the amount of data used to train the model for MP. We use 100%, 50%, 25%, 12%, 6%, and 3% of the QNLI dataset, the largest of the three with 105k examples (3% $\approx$ 3.1k examples), to evaluate the impact of data size on the MP performance. The results are shown in Figure 4 and 5.
Figure 4: MP precision versus flagged examples ($M$) for fractions (3%-100%) of QNLI dataset.
Figure 5: MP precision versus the percentage of the QNLI dataset used for different values of $M$.
We observe a decrease in the MP precision as the size of the dataset decreases in Figure 4. We mentioned earlier that $p_{\theta}(y|x)$ can be approximated as the output probability of a model trained with the leave-one-out dataset when the sample size is large enough. With a decrease in data size, each example plays a more significant role in the learning process. We confirm this as the drop in precision from 6% to 3% is much bigger than drop from 100% to 50%.
In Figure 4, we plot the precision as a function of data size with a fixed $M$. For each $M$, the precision increases as the data size increases but with diminishing returns. In practice, the observed precision can suggest the data size increase required to improve MP performance.
Figure 6: Improvement in model performance (Accuracy/F1-Score) with iterations of simulated human verification using RLC, ALC, and DALC along with MP Precision and the Oracle performance.
@@ 4.5 Iterative Data Correction
The quality of training data improves after each iteration of label correction and the iterative process can be continued until we obtain a high-quality dataset. We simulate correction, without actual human annotators, by using the ground truth known from the original dataset to update the annotations of the flagged data. We fix $M=0.025$ (2.5%) for ATIS and QNLI tasks, but $M=0.050$ (5.0%) for CoNLL because $>50\%$ examples contain noisy tokens. Further, for CoNLL task, we report the F1 score on token-level classification as a standard practice. However, we predict misannotations on the sentence level since the human annotators verify complete sentences containing many tokens. Therefore, we calculate misannotation precision by treating every token as flagged if a sentence is predicted to be misannotated. While this leads to seemingly low precision, it is more useful as a trend compared to sentence-level precision which is consistently high because most examples for CoNLL contain misannotated tokens.
@@
We train models with the updated datasets and plot performance i.e. Accuracy or F1 (using $\times$ markers) after each iteration in Figure 6. We also plot the MP precision for each iteration (using $\bullet$ markers). We say that a model has achieved close-to-oracle performance if the accuracy or F1 score is within a 1% range of the model trained on the original dataset (Oracle). The horizontal green bar in each plot shows this 1% range. For the ATIS task, DALC achieves close-to-oracle performance in 11 iterations. With $M=2.5\%$, this means that the human annotator examines $2.5\%\times 11=27.5\%$ of the dataset which is less than the $29.8\%$ of data misannotated by ChatGPT. ALC takes 2 more iterations i.e. another $2.5\%\times 2=5.0\%$ of data needs correction. For CoNLL, we again observe that DALC requires 11 iterations to achieve close-to-oracle performance. Since $57.4\%$ of CoNLL annotations by ChatGPT contain tokens with incorrect classes, it is reasonable for human annotators to observe $5.0\%\times 11=55\%~{}(<57.4\%)$ of the data. ALC requires another 2 iterations or $10\%$ data-equivalent. For QNLI, we observe close-to-oracle performance after 6 iterations using DALC i.e. correcting $2.5\%\times 6=15\%~{}(<15.1\%)$ of the dataset. ALC requires another iteration or $2.5\%$ data-equivalent. For RLC, the MP precision remains constant as examples are randomly selected for correction. It leads to a much slower improvement in model performance compared to other methods.
@@
We also observe a gradual decrease in MP precision for ALC and DALC due to a decrease in the number of misannotated examples after each iteration. As human annotators examine the flagged examples in practice, we can calculate the MP precision after each iteration and use it as a stopping criterion or as a signal to decrease $M$. Additionally, we hold out a random human-annotated test set to evaluate the model performance after each iteration. If the MP precision is no longer in a target range or the model performance saturates, one can stop performing iterations until more data becomes available from the deployed AI agent.
Overall, we found that DALC can achieve close-to-oracle performance by correcting only a fraction of the dataset. We find that this fraction is less than the fraction of misannotated examples for all three ChatGPT-annotated datasets.

### v3

@@ 3.4 Iterative Data Correction
For each iteration of ALC3, we first train $p_{\theta}$, and perform auto-correction using $p_{\theta}(y^{*}|x)$ as discussed earlier. This is followed by misannotation prediction using $p_{\theta}(y|x)$ where examples are flagged for human annotation. Once the flagged examples are annotated by a human expert, some remaining examples are filtered out to further improve data quality based on low $p_{\theta}(y|x)$. Finally, after above three updates, this dataset is used to train a new model $p_{\theta^{\prime}}$ for the next iteration. In each iteration, a new set of examples are auto-corrected, human-annotated, and filtered. We reset the previously auto-corrected or filtered examples and select a new set based on $p_{\theta^{\prime}}$ predictions. An human-annotated example is not auto-corrected or flagged or filtered in subsequent iterations. The process is repeated iteratively improve the quality of the data and trained model until a stopping criterion (discussed in Section 4.5) is met.
With each iteration, new feedback from the human annotator is accommodated into the dataset and trained model. The new model predictions reveal new inconsistencies with the noisy annotations and lead to the next set of flagged examples. As the count of misannotated examples reduces in the noisy dataset with each iteration, we can expect the MP performance to drop. We discuss this phenomenon in the results section. We can control the MP performance by varying $M$, the fraction of flagged examples. Ideally, one can train a new model after correcting every example but this extreme case is expensive and infeasible due to long waiting time as model training runs after every example. But we still wish to choose $M$ to maximize MP precision and human efficiency. Therefore, in practice, the MP precision can be measured for the latest batch of examples and $M$ can be decreased if precision is low or vice versa. Increasing precision by lowering $M$ comes at the cost of low recall and running more iterations. In our experiments, we use a fixed value of $M$ i.e. a fixed number of examples are flagged in each iteration.
@@ 4.1 Performance on Evaluation Tasks
We evaluate the ALC-based pipeline on three different tasks: (1) ATIS (Airline Travel Information System) Intent Classification Hemphill et al. (1990), (2) CoNLL 2003 Named Entity Recognition or NER Tjong et al. (2003) and (3) Question-answering Natural Language Inference or QNLI Wang et al. (2018). The three tasks are common to many conversational AI agents Lin et al. (2023). Intent classification is used for processing inputs down to different path, NER is used for slot-filling (such as train booking) to fill various parts of a database query, and NLI is used in conversational AI agents to confirm the validity of the output in answering the user query. The three datasets span different levels of complexity and sizes as shown in Table 1. The ground truth labels allow us to compare our pipeline to the oracle performance.
For each task, we created a prompt on OpenAI Playground and used the OpenAI API to get outputs for both train and test splits of each dataset using ‘gpt-3.5-turbo-0613’ model (Sept’21 cut-off) with zero temperature. We use GPT-3.5 as it is available as an API service and performs well at following instructions. For fine-tuning, we used ‘distilbert-base-uncased’, ‘roberta-base’ and ‘albert-large-v2’ pre-trained transformer models from HuggingFace for ATIS, CoNLL and QNLI tasks respectively based on their public leaderboards [^1] https://paperswithcode.com/ and availability on HuggingFace. The results for each task are shown in FT (fine-tuned) column in Table 1. The training details are given in §B.
@@
Table 1: Performance of GPT-3.5 on train and test set of three tasks: (1) ATIS Intent Classification, (2) CoNLL 2003 Name Entity Recognition, and (3) QNLI Natural Language Inference along with performance of fine-tuned models in FT column.
ATIS dataset consists of 17 unique intents (‘airport’, ‘aircraft’, ‘distance’, etc.) assigned to user queries. A description of all 17 intents was provided to GPT-3.5 for classification. 40 examples with multiple labels were removed to reduce the task to single-label classification. For the CoNLL dataset, the prompt describes a sample NER output and four types of entities i.e. person, location, organization, and miscellaneous. QNLI dataset consists of pairs of questions and sentences based on the Stanford Question Answering Dataset Wang et al. (2018). The prompt contains these two inputs and asks if the passage answers the question. All prompts and output parsing are detailed in §C.
The performance of GPT-3.5 is shown in Train and Eval columns in Table 1. The results agree with Qin et al. (2023) in terms of the difference in performance between a fine-tuned model and zero-shot learning using GPT-3. We note that GPT-3.5 has 10-30% lower performance than fine-tuned models across the three tasks. While GPT-3.5 generalizes well to new tasks as a zero/few-shot learner, this result suggests that there is a large performance gap that needs to be bridged when bootstrapping modular AI systems with GPT-3.5. Therefore, it is important to develop a method to improve each module and the overall system over time after deploying the LLM-based AI systems.
4.2 Noise Characteristics
We compare different types of synthetic noise models with the noise induced by GPT-3.5 while annotating the ATIS dataset. We use ATIS dataset for this experiment as it has the highest number of classes which can lead to more pronounced noise patterns. We add different types of synthetic noise to the original dataset with the same proportion of misannotations as GPT-3.5. Table 2 shows the accuracy of models trained with different noises in ATIS dataset and the class imbalance measured by the KL divergence of label distribution from the original label distribution. Details about noise models are provided in Appendix E.
@@ 4.2 Noise Characteristics
GPT-3.5 84.0% 0.0949
Table 2: Comparison of different noise types, including no noise (None), Random, Label-conditional, Input-conditional, and GPT-3.5 annotations, on model Accuracy and Class Imbalance for ATIS dataset.
We find that GPT-3.5 adds the most detrimental noise while having a similar class imbalance as other noise models. We posit that GPT-3.5 provides very reasonable false annotations that are inherently hard to detect or ignore which leads to poor test performance. It is also interesting to note that a model trained on GPT-3.5-annotated training data performs much better than GPT-3.5 itself (84.0% versus 75.6%). This is due to the input data distribution learned by the model leading to better input representations compared to a generic pre-trained LLM. This motivates the fine-tuning of task-specific models for modular AI systems as well as the use of ALC to improve data quality which can lead to further improvements.
In Figure 3, we show 2D t-SNE projections Maaten and Hinton (2008) of Contriever text embeddings Izacard et al. (2022) from ATIS dataset. We visualize examples labeled with 7 classes shown in the legend and larger dots indicate the examples misannotated by GPT-3.5. We observe that most misannotated examples lie near intersections or between class clusters. This indicates that noise induced by GPT-3.5 is indeed complex and depends on the text, its true label, and other possible labels. Previous work has used synthetic random Kremer et al. (2018) or label-conditional noise Rebbapragada et al. (2012); Li et al. (2022) which, as observed in Table 2, are not representative of noise present in GPT-3.5 annotations that we study.
Figure 3: A 2D projection of ATIS text embeddings for a subset of 7 classes. GPT-3.5 annotations are indicated by colors while large dots indicate errors. Most misannotated examples lie near cluster boundaries.
@@ 4.3 Misannotation Prediction
In Figure 4, we compare RLC, ALC, and DALC for MP on all three datasets annotated using GPT-3.5. Note that filtering in ALC3 doesn’t impact misannotation prediction in the first iteration, leading to same results as DALC. We use $\delta=0.75,0.98,0.90$ for ATIS, CoNLL and QLNI to keep the proportion of auto-corrected examples between 3% to 6%. We note from Table 1 that 29.8% of the ATIS and 15.1% of the QNLI train dataset were misannotated by GPT-3.5. For CoNLL, 57.4% of examples contain errors in token classes though only 9.7% of tokens in the dataset are misclassified.
Figure 4: MP precision and recall for ATIS, CoNLL, and QNLI with change in $M$. ALC3 results are same as DALC, and both perform worse than ALC because data quality is improved with auto-correction before MP.
We observe that precision decreases for ALC and DALC as $M$ increases for all tasks because the examples with the highest probability of being misannotated based on $m(.)$ values indeed have a higher rate of being misannotated. In other words, beginning from examples with the highest $m(.)$ values, a human annotator will see a decrease in MP quality over time as they verify examples with lower $m(.)$ values. We also see an increase in recall as $M$ increases because more misannotated examples are flagged at higher values of $M$. We also note that DALC has consistently lower precision and recall than ALC. The auto-corrected examples in DALC typically have high $m(.)$ as $p_{\theta}(y\neq y^{*}|x)$ is low but these examples cannot be flagged. ALC flags the examples with the highest $m(.)$ values leading to higher precision and recall. For RLC, the precision is constant and low while recall increases perfectly linearly with $M$ because examples are flagged at random.
For the human annotator, examples flagged by DALC/ALC3 are more likely to be correct as compared to those flagged by ALC. A more balanced set of flagged examples for the correction task may also avoid the problem of over-reliance of human annotators on model predictions. Further, increase in quality per iteration can be higher with DALC/ALC3 because of two sources of true labels viz. auto-flipping and human corrections. Given a method, we wish to maximize precision to correct a higher proportion of misannotated examples in each iteration. Therefore, we use low values of $M$ in each step in iterative data correction.
@@ 4.4 Effect of Data Size on MP
We wish to understand how the quantity of data collected by modules within an AI system will impact MP performance. Towards this end, we vary the amount of data used to train the model for MP. We use 100%, 50%, 25%, 12%, 6%, and 3% of the QNLI dataset, the largest of the three with 105k examples (3% $\approx$ 3.1k), to evaluate the impact of data size. The results are shown in Figure 5.
(a) MP precision versus flagged examples ($M$) Refer to caption (b) MP precision versus data size (%)
Figure 5: Effect of data size on MP precision. MP precision reduces as more examples are flagged and reduces as data size is decreased, but we observe diminishing returns with increase in data size.
We observe a decrease in the MP precision as the data size decreases (Figure 5(a)). The drop in precision from 6% to 3% is much bigger than drop from 100% to 50% as each example plays a bigger role in learning for small data sizes. In Figure 5(b), we plot the precision as a function of data size with a fixed $M$. The precision increases as the data size increases, but with diminishing returns. In practice, the observed precision can suggest the data size increase required to improve MP performance.
Figure 6: Model performance (Accuracy/F1-score) with iterations of simulated human verification using RLC, ALC, DALC, and ALC3 along with MP Precision and the Oracle performance. Accuracy/F1-score ($\times$) increases while MP precision ($\bullet$) decreases with each iteration. We iterate until close-to-oracle performance is achieved i.e. Accuracy/F1 is within 1% of the ground truth fine-tuned model (green bar).
@@ 4.5 Iterative Data Correction
The quality of training data improves after each iteration of label correction and the iterative process can be continued until we obtain a high-quality dataset. We simulate correction, without actual human annotators, by using ground truth from the original dataset to update the annotations of the flagged examples. We fix $M=0.025$ (2.5%) for ATIS and QNLI tasks, but $M=0.050$ (5.0%) for CoNLL because $>50\%$ examples contain noisy tokens. For CoNLL task, we report the F1-score on token-level classification as a standard practice. However, we predict misannotations on the sentence level since the human annotators verify complete sentences. Therefore, we calculate MP precision by treating every token as flagged if a sentence is predicted to be misannotated. While this artificially leads to low precision, it is more useful to observe compared to sentence-level precision which is consistently high because most examples for CoNLL contain misannotated tokens.
@@
We train models with the updated datasets and plot performance i.e. Accuracy or F1-score ($\times$ markers) after each iteration in Figure 6. We also plot the MP precision ($\bullet$ markers) for each iteration. We say that a model has achieved close-to-oracle performance if the accuracy or F1-score is within a 1% range of the model trained on the original dataset (Oracle). The horizontal green bar in each plot shows this 1% range. For the ATIS task, ALC3, DALC and ALC achieve close-to-oracle performance in 9, 11 and 13 iterations respectively. With $M=2.5\%$, this means only $2.5\%\times 9=22.5\%$ of the dataset is human-annotated with ALC3 which is less than the $29.8\%$ of data misannotated by GPT-3.5. ALC3 requires about 18% less annotations than DALC and 31% less than ALC. For CoNLL, we again observe that ALC3, DALC, and ALC require 9, 11, and 13 iterations to achieve close-to-oracle performance. Since $57.4\%$ of CoNLL annotations by GPT-3.5 contain tokens with incorrect classes, it is reasonable to human-annotate $5.0\%\times 9=45\%~{}(<57.4\%)$ of the data. For QNLI, we need 5 iterations using ALC3 i.e. correcting $2.5\%\times 5=12.5\%~{}(<15.1\%)$ of the dataset. For RLC, MP precision remains constant as examples are randomly selected for correction. It leads to slower improvements in performance compared to other methods. With perfect MP, we would require 29.8%, 57.4%, and 15.1% examples for ATIS, CoNLL, and QNLI respectively to lead to the original dataset quality. But ALC3 has close-to-oracle performance with 17-24% fewer examples than perfect MP by using human feedback in addition to auto-correction and filtering.
@@
We also observe a gradual decrease in MP precision for ALC, DALC, and ALC3 due to a decrease in the number of misannotated examples after each iteration. As human annotators examine the flagged examples in practice, we can calculate the MP precision after each iteration and use it as a stopping criterion or as a signal to decrease $M$. Additionally, as mentioned earlier, we hold out a human-annotated random test set to evaluate model performance. If the MP precision is no longer in a target range or the model performance saturates, one can stop performing iterations until more data becomes available from the deployed AI system. Overall, we find that ALC3 can achieve close-to-oracle performance by correcting only a fraction of the dataset, 17-24% than the fraction of misannotated examples for all three GPT-3.5-annotated datasets, and $\approx 16\%$ less than DALC.